OpenStudy (anonymous):

There is a five-digit number A. With a 1 after it, it is three times as large as with a 1 before it. What is the number?

OpenStudy (anonymous):

Hint: take your original number A and conceive of it as 10000b + 1000c + 100d + 10e + f. You can put a 1 after it and use as the new total number: 100000b + 10000c + 1000d + 100e + 10f + 1. You can then set up equations.

hartnn (hartnn):

i would do something like this 1A = 100000+A A1=A*10 +1 so A*10 +1 = 3(100000+A) fin A from here.......

OpenStudy (anonymous):

oh yeah is a eaual to 42857

hartnn (hartnn):

yup.

OpenStudy (anonymous):

oh thanx mannnn

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

man i need help with one thing

OpenStudy (anonymous):

that i got a 2 madel but don't even increase my point what happen

hartnn (hartnn):

ok, your point will increase only if the medal is given by the person who has asked the question....

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

and its not linear....that 1 medal = 1 point increase....maybe sometimes 2 or more medal = 1point.

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

see., now it increased to 56 ....

OpenStudy (anonymous):

it already

OpenStudy (anonymous):

increase

OpenStudy (anonymous):

not now dude

hartnn (hartnn):

okkk...